# F-91191 Gif-sur-Yvette cedex, France

### v1

@@ 1 Introduction
$N\cdot L=1\ ,\quad\varepsilon_{+}\cdot\varepsilon_{-}=-1\ ,\quad\varepsilon_{+}\cdot N=\varepsilon_{+}\cdot L=0\ ,\quad\eta^{\mu\nu}=2N^{(\mu}L^{\nu)}-2\varepsilon_{+}^{(\mu}\varepsilon_{-}^{\nu)}\ .$ (2)
The peeling theorem, which was initially derived under stronger assumptions than asymptotic simplicity Sachs:1961zz ; Penrose:1962ij , implies that
@@
where $|\vec{x}|$ is the affine parameter along a null geodesics Sachs:1961zz ; Penrose:1972ea . A failure of the peeling property serves as a clear diagnostic for the breakdown of asymptotic simplicity. Since we are interested in identifying the leading terms in the asymptotic expansion of the components of the Weyl tensor, we will only consider the tetrad to leading order in $|\vec{x}|^{-1}$, i.e. at zeroth order. We will choose
@@
$x^{\mu}=t\,t^{\mu}+|\vec{x}|r^{\mu}\ ,\quad\text{with}\quad t^{2}=1,\quad r^{2}=-1\ ,$ (5)
@@
The recent detection of gravitational waves by the LIGO and Virgo collaborations LIGOScientific:2016aoc ; LIGOScientific:2017vwq has catalyzed the development of novel quantum field theory approaches to the classical relativistic two-body problem in general relativity. Significant advances have been made in computing classical scattering observables, such as the scattering angle Bern:2021yeh ; Bern:2021dqo ; Dlapa:2021npj ; Dlapa:2021vgp ; Dlapa:2022lmu ; Jakobsen:2023ndj ; Jakobsen:2023hig ; Damgaard:2023ttc and gravitational waveform Cristofoli:2021vyo ; Jakobsen:2021smu ; Mougiakakos:2022sic ; Riva:2022fru ; Herderschee:2023fxh ; Georgoudis:2023lgf ; Brandhuber:2023hhy ; Caron-Huot:2023vxl ; Bini:2024rsy ; Bohnenblust:2025gir ; Bohnenblust:2023qmy , i.e. the leading $1/|\vec{x}|$ order in the large-$|\vec{x}|$ expansion of the spacetime metric sourced by a scattering process. These methods also shed light on other aspects of general relativity, such as gravitational memory Strominger:2014pwa and the radiation reaction and the tail effect Bern:2021yeh ; Dlapa:2021npj ; Dlapa:2021vgp ; Jakobsen:2023hig ; Jakobsen:2023ndj ; Jakobsen:2022psy ; Elkhidir:2023dco ; Driesse:2024xad ; Driesse:2024feo , and sparked a renewed understanding of the universal properties of soft gravitational radiation Saha:2019tub ; Sahoo:2021ctw ; Alessio:2024onn . They were also used to (re)construct static metrics throughout space-time, such as the Schwarzschild solution, both perturbatively around the origin and to all orders, see e.g. Refs. Boulware:1968zz ; Duff:1973zz ; Bjerrum-Bohr:2002fji ; Neill:2013wsa ; Jakobsen:2020ksu ; Mougiakakos:2020laz ; DOnofrio:2022cvn ; Mougiakakos:2024nku ; Monteiro:2014cda ; Mougiakakos:2024nku ; Damgaard:2024fqj ; Mougiakakos:2024lif .
@@
$\Psi_{0}\sim|\vec{x}|^{-4}\ .$ (6)
This contribution corresponds to the breakdown of asymptotic simplicity found by Damour and Christodoulou in Refs. Damour:1985cm ; Christodoulou:2002 . [^5] Although Ref. Christodoulou:2002 claims that the two results to be equivalent, this connection is not obvious. Stefano De Angelis thanks Thibault Damour for extensive discussions on this point. While asymptotic simplicity is violated, a weaker condition identified by Christodoulou in Ref. Christodoulou:2002 seems to be satisfied. Christodoulou’s condition qualitatively means that the initial data is sufficiently close to Minkowski space and imposes that
$\lim_{|\vec{x}|\rightarrow\infty}\Psi_{1},\Psi_{0}\sim\mathcal{O}\left(|\vec{x}|^{-7/2}\right)\ .$ (7)
This weaker condition is sufficient to prove the global stability of asymptotically-Minkowski spacetimes christodoulou1993global .
@@ 2 The finite distance metric and off-shell non-analytic terms
As discussed, our goal is to compute the NP scalars for the spacetime corresponding to a two-body scattering process and use them as a diagnostic of asymptotic simplicity. The NP scalars $\Psi_{i}$ are nonlinear functions of the metric. In the classical limit and assuming the point-particle approximation, the one-point function $\langle\Psi_{i}\rangle$ factorizes as into a function of the one-point function of its constituents Cristofoli:2021jas ,
@@ 2.1 The finite distance metric from the final-state graviton one-point function
The two factors on the first line are anti-time-ordered and time-ordered, respectively. The first factor in the summand, $\langle\Psi_{\text{in}}|U^{\dagger}(-\infty,t)U^{\dagger}(t,\infty)|n_{\text{fin}}\rangle=\langle n_{\text{fin}}|U(+\infty,-\infty)|\Psi_{\text{in}}\rangle^{*}$, is the operator definition of the conjugated $S$-matrix element between the initial state $|\Psi_{\text{in}}\rangle$ and the final state $|n_{\text{fin}}\rangle$. The second factor, $\langle n_{\text{fin}}|U(+\infty,t){\hat{h}}_{\mu\nu}(x)U(t,-\infty)|\Psi_{\text{in}}\rangle$, is the operator definition of the form factor of the operator ${\hat{h}}_{\mu\nu}(x)$. The leg corresponding to $\hat{h}_{\mu\nu}$ in the first term is not amputated and has a retarded $i\epsilon$ prescription. Importantly, the second term is not identically zero; rather, it subtracts the constant contribution coming from the metric of the two objects when they are very far from each other. [^8] The graviton one-point function can also be defined in the worldline formalism with the retarded propagators of Refs. Mogull:2020sak ; Jakobsen:2021smu . It would be interesting to explore the equivalence of these definitions, perhaps along the lines of Ref. Capatti:2024bid .
@@
$\begin{split}h^{\mu\nu}(x)&=-\int\!\frac{d^{d}k}{(2\pi)^{d}}\frac{e^{-ikx}}{(k_{0}+i\epsilon)^{2}-\vec{k}^{2}}\mathcal{T}^{\mu\nu}_{\mu^{\prime}\nu^{\prime}}J^{\mu^{\prime}\nu^{\prime}}(k)\equiv-\int\!\frac{d^{d}k}{(2\pi)^{d}}\frac{e^{-ikx}{\widetilde{h}}^{\mu\nu}(k)}{(k_{0}+i\epsilon)^{2}-\vec{k}^{2}}\end{split}$ (22)
@@
$\mathcal{T}$ in Eqs. (22) and (23) is the numerator of the graviton propagator, $\zeta$ is an arbitrary (reference) vector and $J^{\mu\nu}(k)$ denotes the classical source generated by the scattering. ${\widetilde{h}}^{\mu\nu}(k)$ is not the direct Fourier transform of the metric due to the presence of the retarded propagator. While this is a non-standard numerator for the graviton propagator, it has the feature that it is proportional to $k^{2}$ when either traced or contracted with the graviton momentum. This feature will be useful in Sec. 4.4 when identifying the leading non-analyticities as $k^{2}\rightarrow 0$.
@@ 2.2 Off-shell and on-shell non-analyticity of the final-state graviton one-point function
The LSZ reduction guarantees that the graviton one-point function differs from its on-shell version by terms that vanish on-shell (i.e. that are proportional to $k^{2}$ and $\varepsilon_{\mu\nu}(k)k^{\mu}$) and by terms connected to the relation between the on-shell and off-shell regulators. In later sections, we will demonstrate in examples and argue in general that additional factors of $k^{2}$ and $\varepsilon_{\mu\nu}(k)k^{\mu}$ compared to the on-shell amplitude lead only to subleading terms in the asymptotic expansion of the position-space scattering metric and that such terms exhibit Sachs’s peeling property. Thus, with the benefit of hindsight, we will ignore these terms and focus on the regularization-related ones. It is moreover natural to expect that the terms that are non-analytic in the components of the graviton momentum behave differently under the Fourier transform (22) than the analytic ones; we will therefore separate the former below.
An off-shell graviton momentum, i.e., $k^{2}\neq 0$, regularizes the IR divergences of S-matrix elements from integrals with three-graviton vertices. These are the integrals that contribute e.g. to the ${\cal O}(G^{3})$ IR divergences of Ref. Weinberg:1965nx and to virtual IR divergences at all loop orders. Thus, on dimensional grounds, we may expect that, on the one hand, there exist terms that are nonanalytic in $k^{2}$ and, on the other, that they can be obtained via a simple map between the dimensional and off-shell regulators.
@@
$I_{1,1,1,0}=\frac{1}{32q_{2}^{2}(\bar{u}_{2}\cdot k)}+\frac{i}{32\pi q_{2}^{2}(\bar{u}_{2}\cdot k)}\left[\ln\frac{(\bar{u}_{2}\cdot k)^{2}}{-k^{2}}+2\ln\frac{q_{2}^{2}}{q_{1}^{2}}\right]+\mathcal{O}\left(\frac{k^{2}}{(\bar{u}_{2}\cdot k)^{2}}\right)\,.$ (25)
where we assumed that $k^{2}<0$. The numerator under the logarithm, $(\bar{u}_{2}\cdot k)^{2}$, is also present in the on-shell calculation, and captures the interaction of the outgoing graviton with its matter source. Comparing with eq. (25) and Eq. (B.5) of Ref. Herderschee:2023fxh we see that they are related, as expected, by the simple replacement
$\frac{1}{\epsilon}-\gamma_{E}+\ln(4\pi^{2})\equiv\frac{1}{\epsilon_{\text{IR}}}~\longmapsto~\ln 4\pi\frac{-k^{2}}{\mu^{2}}\ ,$ (26)
where $\mu$ is the dimensional regularization scale. The same holds for the other integrals with a three-graviton vertex. Thus, the ${\cal O}(G^{3})$ metric can be written as
${\widetilde{h}}^{\mu\nu}_{\text{1-loop}}=-2iG(m_{1}\,u_{1}\cdot k+m_{2}\,u_{2}\cdot k)\log\left(4\pi\frac{-k^{2}}{\mu^{2}}\right){\widetilde{h}}^{\mu\nu}_{\text{tree}}+\text{(analytic in $k^{2}$)}\ .$ (27)
@@
We note that the off-shellness of the graviton momentum does not regularize all infrared divergences, so a dimensional regulator is still necessary to evaluate the final-state graviton one-point function. Indeed, all the integrals appearing in the cut part of the final-state one-point function are IR divergent, even when the external graviton is off-shell. This is because the vertex containing the external graviton also contains an external matter external line. While these divergences can be regularized by taking the matter particles off shell (or, equivalently, preparing the initial state at a finite distance), this will effectively replace the dimensional regularization scale of these divergences by the virtuality of the matter particles. Such terms will not affect the Fourier transform to position space, Eq. (22), so we will not consider this explicitly here. Thus, the general form of the ${\cal O}(G^{3})$ metric in momentum space, separated into nonanalytic and analytic parts, is
${\widetilde{h}}^{\mu\nu}_{\text{1-loop}}$ $=$ $iG(p_{1}+p_{2})\cdot k\left\{2\ln\left[\frac{(u_{1}\cdot k)(u_{2}\cdot k)}{-k^{2}}\right]\right$.
$\left.+\frac{\sigma(\sigma^{2}-3/2)}{(\sigma^{2}-1)^{3/2}}\left(-\frac{1}{\epsilon_{\rm IR}}+\ln\left[\frac{(u_{1}\cdot k)^{2}(u_{2}\cdot k)^{2}}{\mu^{4}}\right]\right)\right\}{\widetilde{h}}^{\mu\nu}_{\text{tree}}+{\mathring{h}}^{\mu\nu}_{\text{1-loop}}\ ,$
where ${\mathring{h}}^{\mu\nu}_{\text{1-loop}}$ is analytic in all components of the graviton momentum and we will not need its explicit form. [^10] The logarithmic branch cuts in Eq. (2.2) have specific $i\epsilon$ prescriptions, which follow from the explicit calculation of the final-state graviton one-point function. We will restore them in Sec. 4 from consistency with the soft-graviton theorems and general properties of the position-space metric., [^11] In writing this expression, we chose a particular organization of the tail terms. In particular, ${\mathring{h}}^{\mu\nu}$ contains $(p_{1}-p_{2})\cdot k\ln(u_{1}\cdot k/u_{2}\cdot k){\widetilde{h}}^{\mu\nu}_{\text{tree}}$. As expected, there is a surviving $1/\epsilon_{\rm IR}$ divergence that originates from the logarithmic drift of the incoming particles due to the long-distance nature of gravitational interactions. Again, this divergence can be removed by shifting the time coordinate or by keeping the incoming particles off-shell as previously mentioned.
It is not difficult to construct off-shell regularization of the IR divergences at higher-loop orders, and we will do so in Sec. 4.4.2. The higher-loop analog of the second term in Eq. (27) requires a nontrivial calculation which we leave for future work.

### v2

@@ 1 Introduction
$N\cdot L=1\ ,\quad\varepsilon_{+}\cdot\varepsilon_{-}=-1\ ,\quad\varepsilon_{+}\cdot N=\varepsilon_{+}\cdot L=0\ ,\quad g^{\mu\nu}=2N^{(\mu}L^{\nu)}-2\varepsilon_{+}^{(\mu}\varepsilon_{-}^{\nu)}\ .$ (2)
The peeling theorem, which was initially derived under stronger assumptions than asymptotic simplicity Sachs:1961zz ; Penrose:1962ij , implies that [^2] Peeling conditions are distinct from a set of weaker conditions identified by Christodoulou in Ref. christodoulou1993global . The latter conditions, which sufficient to prove the global stability of asymptotically-Minkowski spacetimes: $\lim_{|\vec{x}|\rightarrow\infty}\Psi_{1},\Psi_{0}\sim\mathcal{O}\left(|\vec{x}|^{-7/2}\right)\ .$
@@
where $|\vec{x}|$ is the affine parameter along a null geodesics Sachs:1961zz ; Penrose:1972ea and the limit is taken towards future null infinity. A failure of the peeling property serves as a clear diagnostic for the breakdown of asymptotic simplicity. Since we are interested in identifying the leading terms in the asymptotic expansion of the components of the Weyl tensor, we will only consider the tetrad to leading order in $|\vec{x}|^{-1}$, i.e. at zeroth order. We will choose
@@
$x^{\mu}=t\,t^{\mu}+|\vec{x}|r^{\mu}\ ,\quad\text{with }t^{2}=1\,,\;r^{2}=-1\,,\;\text{and }t\cdot r=0\,$, (5)
@@
The recent detection of gravitational waves by the LIGO and Virgo collaborations LIGOScientific:2016aoc ; LIGOScientific:2017vwq has catalyzed the development of novel quantum field theory approaches to the classical relativistic two-body problem in general relativity. Significant advances have been made in computing classical scattering observables, such as the scattering angle Bern:2021yeh ; Bern:2021dqo ; Dlapa:2021npj ; Dlapa:2021vgp ; Dlapa:2022lmu ; Jakobsen:2023ndj ; Jakobsen:2023hig ; Damgaard:2023ttc and gravitational waveform Cristofoli:2021vyo ; Jakobsen:2021smu ; Mougiakakos:2022sic ; Riva:2022fru ; Herderschee:2023fxh ; Georgoudis:2023lgf ; Brandhuber:2023hhy ; Caron-Huot:2023vxl ; Bini:2024rsy ; Georgoudis:2024pdz ; Bohnenblust:2025gir ; Bohnenblust:2023qmy , i.e. the leading $1/|\vec{x}|$ order in the large-$|\vec{x}|$ expansion of the spacetime metric sourced by a scattering process. These methods also shed light on other aspects of general relativity, such as gravitational memory Strominger:2014pwa and the radiation reaction and the tail effect Bern:2021yeh ; Dlapa:2021npj ; Dlapa:2021vgp ; Jakobsen:2023hig ; Jakobsen:2023ndj ; Jakobsen:2022psy ; Kalin:2022hph ; Elkhidir:2023dco ; Driesse:2024xad ; Driesse:2024feo , and sparked a renewed understanding of the universal properties of soft gravitational radiation Saha:2019tub ; Sahoo:2021ctw ; Alessio:2024onn . They were also used to (re)construct static metrics throughout space-time, such as the Schwarzschild solution, both perturbatively around the origin and to all orders, see e.g. Refs. Boulware:1968zz ; Duff:1973zz ; Bjerrum-Bohr:2002fji ; Neill:2013wsa ; Jakobsen:2020ksu ; Mougiakakos:2020laz ; DOnofrio:2022cvn ; Mougiakakos:2024nku ; Monteiro:2014cda ; Mougiakakos:2024nku ; Damgaard:2024fqj ; Mougiakakos:2024lif .
@@
$\Psi_{0}\sim\frac{1}{|\vec{x}|^{4}}\ ,\qquad\Psi_{1}\sim\frac{\log|\vec{x}|}{|\vec{x}|^{4}}\ .$ (6)
This contribution corresponds to the breakdown of asymptotic simplicity found by Damour and Christodoulou in Refs. Damour:1985cm ; Christodoulou:2002 . [^6] Although Ref. Christodoulou:2002 claims that the two results to be equivalent, this connection is not obvious. We thank Thibault Damour for extensive discussions on this point.
@@
Outline of the paper.
The paper is organized as follows. In Sec. 2, we set up the general framework to compute the finite-distance metric from a QFT-based approach as the one-point function of the graviton field. In particular, we provide a generalization of the KMOC formalism to off-shell expectation values, an analyses of the relevant scales and an explicit map to obtain the needed information for the asymptotic analysis by recycling the on-shell computation. In Sec. 3, we describe the technical approach to the asymptotic expansion of the metric toward future null infinity. We identify two regions contributing to the Fourier transform from momentum to position space, in which the graviton momenta are $k^{\mu}\sim|\vec{x}|^{-1}$ and $k^{\mu}\sim|\vec{x}|^{0}$. We refer to them as the Coulomb and radiation regions, respectively. The relevant computations are illustrated in Secs. 3.1 and 3.2. Further technical details on a convenient gauge choice, and the construction of the Weyl tensor and the NP scalars are given in Sec. 3.3. The main results of the paper are discussed in Sec. 4. First we discuss the Coulomb region contribution to the metric: in Sec. 4.1 we study its leading contribution for generic retarded times, as well as $|u|\sim|\vec{x}|$, reproducing the memory displacement between the waveform at early and late retarded times; while, in Sec. 4.2, we focus on finite retarded times, $|u|\ll|\vec{x}|$, and show that the subleading contribution in the Coulomb region leads to the violation of the peeling previously identified in Refs. Damour:1985cm ; Christodoulou:2002 . The analysis of the radiation region is split into two parts: in Sec. 4.3 we focus contributions that are analytic at $k^{2}=0$, and in Sec. 4.4 we analyze the non-analytic contributions, due to long-range non-linear interactions, i.e. the tails. In both cases, we analyze the contributions from the corrections to the tetrad, the non-linearities of the NP scalars in the metric, as well as from singularities at finite distance in the complex $k^{\mu}$ plane. In particular, in Sec. 4.3.3, we generically proof any term which is analytic at $k^{2}=0$ must respect the peeling property. While, in Sec. 4.4.1, combining dimensional analysis with the angular dependence of the Weyl tensor in momentum space, we show that the tails lead to a more severe violation of the peeling property than previously recognized. Finally, we conclude in Sec. 5, raising some open questions about the tension of our new results with the existing results in the literature, and proving a list of future directions.
@@ 2 The finite distance metric and off-shell non-analytic terms
As discussed, our goal is to compute the NP scalars for the spacetime corresponding to a two-body scattering process and use them as a diagnostic of asymptotic simplicity. The NP scalars $\Psi_{i}$ are nonlinear functions of the metric. In the classical limit and assuming the point-particle approximation, the one-point function $\langle\Psi_{i}\rangle$ factorizes into a function of the one-point function of its constituents Cristofoli:2021jas ,
@@ 2.1 The finite distance metric from the final-state graviton one-point function
The two factors on the first line are anti-time-ordered and time-ordered, respectively. The first factor in the summand, $\langle\Psi_{\text{in}}|U^{\dagger}(-\infty,t)U^{\dagger}(t,\infty)|n_{\text{fin}}\rangle=\langle n_{\text{fin}}|U(+\infty,-\infty)|\Psi_{\text{in}}\rangle^{*}$, is the operator definition of the conjugated $S$-matrix element between the initial state $|\Psi_{\text{in}}\rangle$ and the final state $|n_{\text{fin}}\rangle$. The second factor, $\langle n_{\text{fin}}|U(+\infty,t){\hat{h}}_{\mu\nu}(x)U(t,-\infty)|\Psi_{\text{in}}\rangle$, is the operator definition of the form factor of the operator ${\hat{h}}_{\mu\nu}(x)$. The leg corresponding to $\hat{h}_{\mu\nu}$ in the first term is not amputated and has a retarded $i\epsilon$ prescription. Importantly, the second term is not identically zero; rather, it subtracts the constant contribution coming from the metric of the two objects when they are very far from each other. [^9] The graviton one-point function can also be defined in the worldline formalism with the retarded propagators of Refs. Kalin:2020mvi ; Mogull:2020sak ; Jakobsen:2021smu . It would be interesting to explore the equivalence of these definitions, perhaps along the lines of Ref. Capatti:2024bid .
@@
$\begin{split}h^{\mu\nu}(x)&=-\int\!\frac{d^{d}k}{(2\pi)^{d}}\frac{e^{-ik\cdot x}}{(k_{0}+i\epsilon)^{2}-\vec{k}^{2}}\mathcal{T}^{\mu\nu}_{\mu^{\prime}\nu^{\prime}}J^{\mu^{\prime}\nu^{\prime}}(k)\equiv-\int\!\frac{d^{d}k}{(2\pi)^{d}}\frac{e^{-ik\cdot x}{\widetilde{h}}^{\mu\nu}(k)}{(k_{0}+i\epsilon)^{2}-\vec{k}^{2}}\end{split}$ (21)
@@
$\mathcal{T}$ in Eqs. (21) and (22) is the numerator of the graviton propagator, $\zeta$ is an arbitrary (reference) vector and $J^{\mu\nu}(k)$ denotes the classical source generated by the scattering, including the matter stress-tensor and the gravitational non-linearities. ${\widetilde{h}}^{\mu\nu}(k)$ is not the direct Fourier transform of the metric due to the presence of the retarded propagator. While this is a non-standard numerator for the graviton propagator, it has the feature that it is proportional to $k^{2}$ when either traced or contracted with the graviton momentum. This feature will be useful in Sec. 4.4 when identifying the leading non-analyticities as $k^{2}\rightarrow 0$.
@@ 2.2 Off-shell and on-shell non-analyticity of the final-state graviton one-point function
Scattering amplitudes are typically infrared divergent even in the classical limit. Dimensional regularization, defining then in $d=4-2\epsilon$ dimensions, provides a gauge-invariant regularization scheme. Since we are interested in computing the metric at finite distance, the momentum associated to the graviton is to be considered off-shell, i.e. $k^{2}\neq 0$. While this off-shellness regularizes the infrared divergences associated with the outgoing external graviton, it is not a substitute of dimensional regularization which is still required to regularize certain certain initial-state singularities. In this section we describe a systematic procedure to obtain the leading logarithms in $k^{2}$ of the relevant integrals from the dimensionally-regulated ones.
The LSZ reduction guarantees that the graviton one-point function differs from its on-shell version by terms that vanish on-shell (i.e. that are proportional to $k^{2}$ and $\varepsilon_{\mu\nu}(k)k^{\mu}$) and by terms connected to the relation between the on-shell and off-shell regulators. In later sections, we will demonstrate in examples and argue in general that additional factors of $k^{2}$ and $\varepsilon_{\mu\nu}(k)k^{\mu}$ compared to the on-shell amplitude lead only to subleading terms in the asymptotic expansion of the position-space scattering metric and that such terms exhibit Sachs’ peeling property. Thus, with the benefit of hindsight, we will ignore these terms and focus on the regularization-related ones. It is moreover natural to expect that the terms that are non-analytic in the components of the graviton momentum behave differently under the Fourier transform (21) than the analytic ones; we will therefore separate the former below.
An off-shell graviton momentum, i.e., $k^{2}\neq 0$, regularizes the IR divergences of correlation functions from diagrams with three-graviton vertices. These are the integrals that contribute e.g. to the ${\cal O}(G^{3})$ IR divergences of Ref. Weinberg:1965nx and to virtual IR divergences at all loop orders. Thus, on dimensional grounds, we may expect that, on the one hand, there exist terms that are nonanalytic in $k^{2}$ and, on the other, that they can be obtained via a simple map between the dimensional and off-shell regulators.
@@
$I_{1,1,1,0}$ $=\int\frac{\textrm{d}^{4}\ell}{(2\pi)^{4}}\frac{\hat{\delta}(2u_{2}\cdot\ell)}{\ell^{2}(\ell+q_{2})^{2}(\ell-q_{1})^{2}}$
$=\frac{1}{16q_{2}^{2}(u_{2}\cdot k)}+\frac{i}{16\pi q_{2}^{2}(u_{2}\cdot k)}\log\frac{4(u_{2}\cdot k)^{2}q_{2}^{2}}{(-k^{2})q_{1}^{2}}+\mathcal{O}\left(\frac{k^{2}}{(u_{2}\cdot k)^{2}}\right)\,.$ (24)
where we have assumed that $k^{2}<0$. The numerator under the logarithm, $(u_{2}\cdot k)^{2}$, is also present in the on-shell calculation, and captures the interaction of the outgoing graviton with its matter source. The same integral has been computed with dimensional regularization and $k^{2}=0$ for the computation of the one-loop waveform, see e.g. Eq. (B.5) of Ref. Herderschee:2023fxh . Comparing with eq. (24), we see that the regulator dependent part are related, as expected, by the simple replacement
$\frac{1}{\epsilon}-\gamma_{E}+\log(4\pi^{2})\equiv\frac{1}{\epsilon_{\text{IR}}}~\longmapsto~2\log\frac{\pi(-k^{2})}{\mu^{2}}\ ,$ (25)
where $\mu$ is the dimensional regularization scale. [^11] The real part of eq. (24) turns out to differ from the same result in dimensional regularization computed in Refs. Brandhuber:2023hhy ; Herderschee:2023fxh (the rational part differ by a factor of two, as well as the relative coefficient between the different logarithms). Despite this apparent difference, we believe that the two regularizations will still result in the same waveform up to a different, and scheme-dependent, phase. Indeed, we note that waveform result will look very different from the one computed within dimensional regularization and on shell: with the current $k^{2}\neq 0$ regularization, the master integral basis for the waveform will be different due to this extra scale, and so will be their coefficients due to different IBP identities. We expect this to reconcile the difference in individual master integrals. An explicit confirmation is of course desirable, but it is beyond the scope of the present paper. The same holds for the other integrals with a three-graviton vertex. Thus, the ${\cal O}(G^{3})$ metric can be written as
${\widetilde{h}}^{\mu\nu}_{\text{1-loop}}=-2iG(m_{1}\,u_{1}+m_{2}\,u_{2})\cdot k\,\log\left[\frac{\pi(-k^{2})}{\mu^{2}}\right]{\widetilde{h}}^{\mu\nu}_{\text{tree}}+\text{(analytic in $k^{2}$)}\ .$ (26)
@@
We note that the off-shellness of the graviton momentum does not regularize all infrared divergences, so a dimensional regulator is still necessary to evaluate the final-state graviton one-point function. Indeed, the propagators on the matter lines appear with a retarded $i\epsilon$ prescription. While the time-symmetric part of these contributions does not need extra regulators in the off-shell one-point function, the delta function introduces extra IR divergences associated with the incoming matter states Caron-Huot:2023vxl . While these divergences can be regularized by taking the matter particles off shell (or, equivalently, preparing the initial state at a finite distance), this will effectively replace the dimensional regularization scale of these divergences by the virtuality of the matter particles. Such terms will not affect the Fourier transform to position space, Eq. (21), so we will not consider this explicitly here. In particular, the surviving $1/\epsilon_{\rm IR}$ divergence originates from the logarithmic drift of the incoming particles due to the long-distance nature of gravitational interactions. [^12] This divergence can be removed by shifting the time coordinate or by keeping the incoming particles off-shell as previously mentioned. In the latter case, the external propagators must have advanced $i\epsilon$ prescriptions, as they correspond to the boundary conditions in the far past.
Thus, the general form of the ${\cal O}(G^{3})$ metric in momentum space, separated into nonanalytic and analytic parts, is
${\widetilde{h}}^{\mu\nu}_{\text{1-loop}}$ $=$ $iG(m_{1}u_{1}+m_{2}u_{2})\cdot k\left\{2\log\left[\frac{(u_{1}\cdot k)(u_{2}\cdot k)}{-k^{2}}\right]\right$.
$\left.+\frac{\sigma(\sigma^{2}-3/2)}{(\sigma^{2}-1)^{3/2}}\left(-\frac{1}{\epsilon_{\rm IR}}+\log\left[\frac{(u_{1}\cdot k)^{2}(u_{2}\cdot k)^{2}}{\mu^{4}}\right]\right)\right\}{\widetilde{h}}^{\mu\nu}_{\text{tree}}+{\mathring{h}}^{\mu\nu}_{\text{1-loop}}\ ,$
where ${\mathring{h}}^{\mu\nu}_{\text{1-loop}}$ is analytic in all components of the graviton momentum and we will not need its explicit form. [^13] The logarithmic branch cuts in Eq. (2.2) have specific $i\epsilon$ prescriptions, which follow from the explicit calculation of the final-state graviton one-point function. We will restore them in Sec. 4 from consistency with the soft-graviton theorems and general properties of the position-space metric., [^14] In writing this expression, we chose a particular organization of the tail terms. In particular, ${\mathring{h}}^{\mu\nu}$ contains $(p_{1}-p_{2})\cdot k\log(u_{1}\cdot k/u_{2}\cdot k){\widetilde{h}}^{\mu\nu}_{\text{tree}}$. Indeed, in Sec. 4.3.3, we will present a proof that analyticity at $k^{2}=0$ implies the peeling conditions.
It is not difficult to construct an off-shell regularization of the IR divergences at higher-loop orders, and we will do so in Sec. 4.4.2. The higher-loop analog of the second term in Eq. (26) requires a nontrivial calculation which we leave for future work.